# Unveiling AI’s Blind Spots: An Oracle for In-Domain, Out-of-Domain, and Adversarial Errors

### v2

@@ Appendix E More comparisons between SuperMentor and baselines
As noted in the caption of Tab. 1, baseline performances across multiple hyperparameter configurations are indicated in Appendix, Tab. S2. In addition, when ViT serves as the mentee, the SuperMentor’s performance is shown in Appendix, Tab. S3. It can be seen that our SuperMentor can still achieve the best performance when the mentee has a backbone of ViT. This emphasizes the importance of the error source used during training.
@@
ConfidNet (Corbière et al., 2019) | 67.3 | 65.9 | 68.5 | 70.8 | 67.9 | 68.8 | 69.0 | 70.1 | 67.1 | 68.4 |
TrustScore (Jiang et al., 2018) | 70.0 | 67.5 | 71.1 | 73.4 | 71.3 | 73.2 | 72.5 | 73.9 | 72.7 | 71.7 |
SSL (Luo et al., 2021) | 66.3 | 66.7 | 69.7 | 71.9 | 70.0 | 69.1 | 70.1 | 70.2 | 66.7 | 69.0 |
SuperMentor (ours) | 78.9 | 73.6 | 74.6 | 78.3 | 73.6 | 83.0 | 78.4 | 79.9 | 72.3 | 77.0 |
@@
ConfidNet (Corbière et al., 2019) | 73.3 | 74.1 | 73.4 | 75.6 | 74.5 | 69.2 | 70.5 | 66.9 | 70.5 | 72.0 |
TrustScore (Jiang et al., 2018) | 73.7 | 72.0 | 72.4 | 73.5 | 72.5 | 70.4 | 70.1 | 69.2 | 72.5 | 71.8 |
SSL (Luo et al., 2021) | 64.3 | 66.3 | 68.2 | 70.6 | 69.3 | 62.5 | 66.0 | 60.1 | 62.4 | 65.5 |
SuperMentor (ours) | 72.7 | 71.3 | 71.7 | 73.0 | 70.2 | 80.2 | 74.1 | 72.9 | 74.1 | 73.4 |
Appendix F Ablation study
As mentioned in Sec. 4.5, we examine the effect of the distillation loss (Fig. 2) on the SuperMentor performance. The results are presented in Appendix, Tab. S4. It is clear that excluding results in a decrease in SuperMentor’s accuracy across all datasets. For example, in the C10 dataset, the average accuracy of SuperMentor decreases from to . This suggests that encourages SuperMentor to learn the fine-grained decision boundaries among different object classes of a mentee.
Alternatively, instead of utilizing the mentee’s logits, SuperMentor can incorporate an additional cross-entropy loss to align the mentor’s predicted object class labels with those of the mentee, denoted as . From Appendix, Tab. S4, we observe that replacing with leads to a slight decrease in accuracy. This is due to the fact that the mentee’s logits contain more information than the mentee’s class labels.
OOD | AA | |||||||||||
ID | SpN | GaB | Spat | Sat | PGD | CW | Jitter | PIFGSM | Average | |||
✗ | ✗ | 57.5± 1.2 | 61.0± 0.8 | 56.1± 1.1 | 58.6± 0.6 | 54.3± 1.5 | 58.6± 0.5 | 59.1± 1.2 | 58.5± 1.0 | 59.6± 0.6 | 58.2± 2.0 | |
✗ | ✓ | 80.0± 1.8 | 73.7± 0.6 | 79.2± 2.0 | 77.9± 0.8 | 74.3± 1.9 | 80.5± 0.8 | 76.5± 0.8 | 79.7± 0.7 | 71.2± 0.4 | 77.0± 3.2 | |
C10 | ours | 80.9± 1.6 | 73.2± 0.7 | 80.5± 1.4 | 79.4± 1.3 | 75.6± 1.0 | 81.4± 0.9 | 78.2± 0.9 | 80.7± 1.2 | 71.9± 0.1 | 78.0± 3.5 | |
✗ | ✗ | 56.8± 1.2 | 59.5± 0.8 | 56.6± 1.1 | 57.8± 1.2 | 53.7± 1.3 | 57.7± 1.9 | 57.3± 1.9 | 57.3± 1.7 | 57.1± 0.5 | 57.1± 1.8 | |
✗ | ✓ | 75.0± 0.7 | 70.9± 0.3 | 74.8± 0.7 | 74.1± 0.3 | 68.1± 0.8 | 78.1± 0.6 | 75.2± 0.6 | 76.2± 0.5 | 66.5± 1.0 | 73.2± 3.7 | |
C100 | ours | 75.4± 0.7 | 71.1± 0.1 | 75.4± 0.6 | 74.5± 0.8 | 68.4± 1.1 | 78.3± 0.4 | 75.6± 0.4 | 76.6± 0.2 | 66.9± 0.4 | 73.6± 3.7 | |
✗ | ✗ | 73.0± 4.2 | 70.1± 2.8 | 69.6± 3.1 | 72.8± 3.7 | 68.5± 3.4 | 75.8± 5.2 | 72.5± 4.2 | 73.6± 4.6 | 70.7± 0.5 | 71.9± 3.8 | |
✗ | ✓ | 78.7± 0.1 | 73.1± 0.2 | 73.6± 0.5 | 78.0± 0.1 | 73.2± 0.3 | 83.0± 0.1 | 78.4± 0.1 | 79.9± 0.1 | 72.2± 0.2 | 76.7± 3.6 | |
IN | ours | 78.9± 0.0 | 73.6± 0.1 | 74.6± 0.2 | 78.3± 0.1 | 73.6± 0.2 | 83.0± 0.1 | 78.4± 0.1 | 79.9± 0.0 | 72.3± 0.1 | 77.0± 3.4 |
@@ Appendix H SuperMentor in real-world practice
To reflect the real-world contribution of our work, as mentioned in Sec. 5, we expanded our experiments to the NCTCRCHE100K (Kather et al., 2019) dataset which is used for medical image classification on colorectal cancer. This dataset comprises around 100K images across 9 tissue classes. Predicting AI models’ errors in medical image classification has significant real-world implications, such as reducing misdiagnoses and increasing the reliability of AI-assisted medical tools. Tab. S5 demonstrates that our SuperMentor accurately predicts the correctness of mentee outputs with an average accuracy of 81.9%, significantly outperforming other baseline methods. This highlights that our proposed framework can offer greater value and reliability for AI error prediction compared to existing approaches in medical image classification.
@@
SER | 49.8 | 49.6 | 50.4 | 50.0 | 49.9 | 50.1 | 50.2 | 49.8 | 49.8 | 50.0 |
MCP () | 53.9 | 59.4 | 54.3 | 54.6 | 50.4 | 54.2 | 54.4 | 54.7 | 55.0 | 54.6 |
MCP () | 65.1 | 71.3 | 65.0 | 66.2 | 49.9 | 66.2 | 66.1 | 65.8 | 67.1 | 64.7 |
MCP () | 74.5 | 78.7 | 74.3 | 76.2 | 49.4 | 76.9 | 75.8 | 75.9 | 76.2 | 73.1 |
CPE () | 78.5 | 74.2 | 81.9 | 77.7 | 45.2 | 77.4 | 77.4 | 77.8 | 77.3 | 74.2 |
CPE () | 77.6 | 79.3 | 77.8 | 78.2 | 48.8 | 79.2 | 78.0 | 78.8 | 78.3 | 75.1 |
CPE () | 67.6 | 76.3 | 67.0 | 68.8 | 51.7 | 70.3 | 69.0 | 69.9 | 69.5 | 67.8 |
DTC () | 77.7 | 67.3 | 77.2 | 79.7 | 66.2 | 78.8 | 77.4 | 77.6 | 78.1 | 75.5 |
DTC () | 69.4 | 48.5 | 67.2 | 61.6 | 52.4 | 59.2 | 64.7 | 62.2 | 59.2 | 60.5 |
DTC () | 52.6 | 49.4 | 53.2 | 50.7 | 50.0 | 51.1 | 52.4 | 51.5 | 50.9 | 51.3 |
ConfidNet (Corbière et al., 2019) | 83.6 | 59.7 | 89.4 | 74.0 | 56.5 | 80.1 | 85.7 | 82.1 | 80.4 | 76.8 |
TrustScore (Jiang et al., 2018) | 73.3 | 62.9 | 75.2 | 63.7 | 63.7 | 73.1 | 74.9 | 73.5 | 74.0 | 70.5 |
SSL (Luo et al., 2021) | 79.3 | 57.1 | 87.4 | 75.9 | 56.9 | 76.2 | 83.1 | 77.7 | 76.9 | 74.5 |
SuperMentor (ours) | 88.7 | 64.9 | 69.8 | 85.0 | 54.5 | 89.4 | 88.7 | 88.2 | 87.4 | 81.9 |
@@ Appendix I SuperMentor performance on additional OOD domains on ImageNet dataset
As mentioned in Sec. 5, we included experiments to evaluate the SuperMentor in more diversified OOD domains on ImageNet datasets, including ImageNet9-MIXED-RAND (IN9-MR) (Xiao et al., 2020), ImageNet9-MIXED-SAME (IN9-MS) (Xiao et al., 2020), ImageNet9-MIXED-NEXT (IN9-MN) (Xiao et al., 2020), ImageNet-R (IN-R) (Hendrycks et al., 2021) and ImageNet-Sketch (IN-S) (Wang et al., 2019). Specifically, the MIXED-RAND, MIXED-SAME, and MIXED-NEXT datasets are derived from 9 classes in ImageNet and contain varying amounts of background and foreground signals. These datasets aim to demonstrate that models often classify objects based on background cues (often spurious features), rather than the objects themselves. Specifically, MIXED-SAME, MIXED-RAND, and MIXED-NEXT represent images with random backgrounds from the same class, random backgrounds from a random class, and random backgrounds from the next class, respectively. The ImageNet-R dataset comprises images featuring artistic representations of objects, such as cartoons, community-generated art, and graffiti renditions. The ImageNet-Sketch dataset consists of sketch-like images that match the ImageNet validation set in both categories and scale.
@@
To achieve a more thorough evaluation of our SuperMentor (see Sec. 4.5) in OOD domains, we assess its performance on these additional OOD datasets. It is worth noting that the SuperMentor learns from the mentee’s errors on adversarial ImageNet images generated by the PIFGSM attack only. In other words, the SuperMentor was NOT trained on any of these additional ImageNet-related OOD datasets. The SuperMentor achieves average accuracies of 68.46%, 69.24%, 67.04%, 58.68%, and 58.68% on the IN9-MR, IN9-MS, IN9-MN, IN-R, and IN-S datasets respectively, surpassing the 50% chance level. Despite that our Supermentor model is a simple ViT, it is still quite remarkable to achieve above-chance error prediction performance. We hope our work inspires researchers to explore more sophisticated AI mentors, incorporating advanced architectures or loss functions, to further enhance the ability to predict the correctness of another AI model’s outputs.
Appendix J Discussion on the vulnerabilities of the mentor AI
Beyond error prediction, our framework improves the understanding of both individual and systemic vulnerabilities in Deep Neural Network (DNN)-based mentees, fostering more robust AI systems. This is further illustrated by the EigenCAM (Gildenblat & contributors, 2021; Muhammad & Yeasin, 2020) visualization in Appendix, Fig. S2. From these visualizations, we can make two observations below:
The vulnerabilities of the mentee do not overlap with those of the mentors. As illustrated in Appendix, Fig. S2, the mentor does not merely replicate the mentee’s learning patterns, as their activation maps for identical input images can differ significantly. For example, in the images shown in Appendix, Fig. S2 (a), although the mentor and mentee exhibit very similar activation maps for ID images, their activations diverge greatly when the input image is corrupted by Gaussian Blur (GaB) or subjected to adversarial attacks using PIFGSM method. In the IN-OOD-GaB and IN-AA-PIFGSM images, the mentee demonstrates vulnerability under both corruption and adversarial attacks, whereas our SuperMentor does not exhibit such vulnerabilities on these OOD and AA images, consistently focusing on the brambling object. This indicates the vulnerabilities of mentors and mentees are distinct due to their non-overlapping objectives.
Our AI mentor serves as a valuable diagnostic tool for AI mentees. We argue that, beyond merely detecting errors, mentors function as important diagnostic tools for AI mentees, as supported in Appendix, Fig. S2. For example, in Appendix, Fig. S2 (b), although the mentee correctly classifies the ID image, its activation map does not concentrate on the baseball but instead highlights background cues (spurious features). When the input image is corrupted by Gaussian Blur (GaB) or subjected to adversarial PIFGSM attacks, the mentee fails to make accurate predictions. In contrast, our SuperMentor effectively focuses on the stitches of the baseball—a key feature—regardless of whether the image is corrupted or adversarially attacked. The mentor highlights the stitches of the baseball as the key feature pattern that the mentee tends to overlook. This suggests that the mentor can identify vulnerabilities in the DNN-based model, even though it shares the same ViT architecture as the mentee. Furthermore, this reveals significant potential for our mentor to help refine the mentee’s performance by providing valuable insights into the mentee’s behavior.
In conclusion, Appendix, Fig. S2 provides strong evidence that the AI mentor not only does not share the same vulnerabilities as the mentee but also serves as a valuable diagnostic tool for AI mentees.
@@
Appendix K Detailed performance of mentors across various error sources
@@
As mentioned in the caption of Fig. 3, the detailed results of mentors across various error sources for the CIFAR-10, CIFAR-100, ImageNet-1K datasets are shown in Appendix, Fig. S3, Fig. S4 and Fig. S5 respectively.
Appendix L Detailed performance of mentors across mentee architectures
In Fig. 5, we show the generalization performance of mentors averaged over all three error types of mentees with various architectures. Here, we expand the results in the form of tables listing out all the individual accuracy for all the error sources on CIFAR-10, CIFAR-100 and ImageNet-1K datasets in Appendix, Tab. S6, Tab. S7, and Tab. S8 respectively.
Mentor | ResNet50 | ViT | |||
Mentee | ResNet50ViT | ViTResNet50 | ResNet50ViT | ViTResNet50 | |
ID | 59.32.859.62.0 | 56.73.654.52.4 | 65.73.365.63.4 | 63.73.862.53.0 | |
SpN | 59.13.158.43.3 | 57.52.857.42.1 | 67.94.466.15.0 | 64.53.562.92.7 | |
GaB | 58.12.258.22.0 | 59.11.657.31.8 | 62.73.162.42.8 | 63.83.262.62.8 | |
Spat | 59.42.158.42.1 | 58.03.457.43.0 | 66.62.866.13.7 | 66.84.565.23.8 | |
OOD | Sat | 57.11.957.71.3 | 58.71.256.71.8 | 65.94.366.43.2 | 68.84.367.74.3 |
PGD | 60.52.258.41.5 | 64.92.762.31.7 | 71.45.569.04.5 | 77.02.672.34.1 | |
CW | 61.62.759.22.2 | 64.13.060.32.0 | 69.54.667.74.0 | 71.42.667.63.1 | |
Jitter | 60.12.659.31.8 | 62.93.860.13.3 | 71.74.270.43.5 | 74.42.470.43.8 | |
AA | PIFGSM | 67.54.165.33.4 | 64.03.062.12.3 | 78.03.575.43.4 | 77.02.672.53.9 |
Average | 60.33.959.43.2 | 60.74.258.73.4 | 68.85.967.75.1 | 69.76.267.15.2 |
Mentor | ResNet50 | ViT | |||
Mentee | ResNet50ViT | ViTResNet50 | ResNet50ViT | ViTResNet50 | |
ID | 58.91.958.41.8 | 61.22.760.12.1 | 61.82.861.42.9 | 65.63.463.82.0 | |
SpN | 64.02.061.83.5 | 57.52.357.92.9 | 68.71.965.1 3.2 | 65.43.264.62.6 | |
GaB | 59.32.658.62.0 | 61.72.961.01.9 | 65.02.263.82.5 | 66.53.564.61.9 | |
Spat | 58.71.657.71.9 | 60.02.259.52.0 | 66.42.864.92.7 | 66.13.364.02.0 | |
OOD | Sat | 59.63.558.83.4 | 57.41.957.42.6 | 65.82.164.43.0 | 58.94.458.44.4 |
PGD | 60.71.359.41.1 | 62.02.360.61.3 | 67.52.665.52.0 | 71.04.067.22.1 | |
CW | 60.52.659.21.8 | 60.82.958.92.1 | 66.43.664.52.1 | 66.33.863.61.8 | |
Jitter | 58.62.457.82.0 | 62.43.760.52.4 | 67.23.465.22.2 | 67.43.765.11.9 | |
AA | PIFGSM | 67.34.664.62.5 | 66.24.164.32.2 | 73.63.769.11.7 | 72.14.468.52.3 |
Average | 60.83.959.63.2 | 61.03.860.02.9 | 66.94.164.93.2 | 66.65.264.43.6 |
Mentor | ResNet50 | ViT | |||
Mentee | ResNet50ViT | VResNet50 | ResNet50ViT | ViTResNet50 | |
ID | 65.13.363.92.7 | 62.21.362.51.6 | 73.32.371.02.4 | 71.21.871.51.4 | |
SpN | 62.32.462.03.7 | 58.81.658.42.2 | 73.91.471.12.2 | 71.42.371.21.3 | |
GaB | 62.72.862.93.8 | 60.72.360.52.7 | 73.92.271.83.0 | 72.02.971.51.7 | |
Spat | 61.81.561.11.7 | 61.91.561.71.4 | 73.42.071.42.3 | 71.82.471.71.4 | |
OOD | Sat | 64.32.863.93.5 | 62.33.062.12.7 | 73.21.971.22.6 | 70.62.470.31.2 |
PGD | 69.35.066.82.6 | 69.04.967.12.7 | 76.63.872.62.0 | 71.33.270.31.6 | |
CW | 66.05.564.33.5 | 67.44.265.82.3 | 74.03.671.11.9 | 73.22.971.71.4 | |
Jitter | 66.94.165.32.5 | 68.04.566.12.6 | 74.93.072.22.0 | 73.33.171.31.3 | |
AA | PIFGSM | 72.34.569.32.2 | 69.44.768.02.9 | 77.03.472.91.7 | 73.32.772.11.4 |
Average | 65.65.064.43.8 | 64.45.163.63.9 | 74.43.071.72.4 | 72.02.871.31.5 |

### v3

@@ Appendix E More comparisons between SuperMentor and baselines
As noted in the caption of Fig. 6(a), baseline performances across multiple hyperparameter configurations are indicated in Appendix, Tab. S1. In addition, when ViT serves as the mentee, the SuperMentor’s performance is shown in Appendix, Tab. S2. It can be seen that our SuperMentor can still achieve the best performance when the mentee has a backbone of ViT. This emphasizes the importance of the error source used during training.
@@
| ConfidNet (Corbière et al., 2019) | 67.8 | 59.0 | 61.1 | 61.8 | 64.1 | 51.8 | 61.9 | 57.5 | 54.3 | 59.3 |
| TrustScore (Jiang et al., 2018) | 71.2 | 58.6 | 60.1 | 63.4 | 65.7 | 58.2 | 63.3 | 59.4 | 60.8 | 61.6 |
| SSL (Luo et al., 2021) | 68.6 | 66.2 | 69.2 | 69.6 | 68.4 | 64.7 | 63.4 | 63.2 | 67.9 | 66.7 |
| SuperMentor (ours) | 73.3 | 69.2 | 65.6 | 71.1 | 69.9 | 73.1 | 67.4 | 70.3 | 75.4 | 70.4 |
@@
| ConfidNet (Corbière et al., 2019) | 73.4 | 69.0 | 67.0 | 68.6 | 71.1 | 58.8 | 59.8 | 46.8 | 60.7 | 63.2 |
| TrustScore (Jiang et al., 2018) | 74.4 | 67.5 | 64.8 | 67.6 | 70.8 | 57.9 | 59.7 | 48.1 | 60.5 | 62.6 |
| SSL (Luo et al., 2021) | 66.7 | 67.2 | 68.0 | 66.1 | 66.9 | 58.1 | 56.7 | 43.4 | 59.2 | 60.9 |
| SuperMentor (ours) | 70.9 | 65.5 | 61.7 | 67.5 | 68.8 | 81.6 | 64.1 | 68.5 | 81.9 | 70.0 |
Appendix F Relationship between mentor generalization and natural adversarial samples
As discussed in Sec. 4.4, we aim to examine the relationship between the generalization ability of mentors and natural adversarial samples. Following the definition in (Hendrycks et al., 2021b), natural adversarial samples are those that consistently lead to incorrect predictions across a wide range of AI models. In line with this definition, for each error source, we define as the set of samples misclassified by both ResNet50 and ViT mentees, serving as an analogue to the natural adversarial samples in (Hendrycks et al., 2021b). The samples misclassified only by the ResNet50 mentee (the ViT mentee can classify them correctly) are labeled set , and those misclassified only by the ViT mentee (the ResNet50 mentee can classify them correctly) are labeled set . The SuperMentor is trained on performance from the ResNet50 mentee and then generalizes its error prediction ability to the ViT mentee.
The result in Appendix, Tab. S3 indicates that the natural adversarial samples in play a key role in explaining SuperMentor’s strong generalization across different architectures. However, SuperMentor can also perform well on the non-natural adversarial samples ( and ). For example, it achieves 76.5% accuracy on the set for ID error types. This demonstrates that SuperMentor’s robust generalization ability is not limited solely to natural adversarial samples.
| Error Source | |||
| ID | 1750/2225 (78.7%) | 1055/1352 (78.0%) | 429/561 (76.5%) |
| SpN | 2568/2973 (86.4%) | 2310/3075 (75.1%) | 462/579 (79.8%) |
| GaB | 2710/3078 (88.0%) | 1368/1657 (82.6%) | 642/757 (84.8%) |
| Spat | 2131/2587 (82.4%) | 1300/1627 (79.9%) | 464/584 (79.5%) |
| Sat | 2433/2880 (84.5%) | 1585/2012 (78.8%) | 502/618 (81.2%) |
| PGD | 3785/4689 (80.7%) | 377/709 (53.2%) | 2360/3286 (71.8%) |
| CW | 2528/3281 (77.0%) | 977/1952 (61.4%) | 1325/1886 (70.3%) |
| Jitter | 2527/3501 (72.2%) | 701/1032 (67.9%) | 1714/2685 (63.8%) |
@@ Appendix H SuperMentor in real-world practice
To reflect the real-world contribution of our work, as mentioned in Sec. 5, we expanded our experiments to the NCTCRCHE100K (Kather et al., 2019) dataset which is used for medical image classification on colorectal cancer. This dataset comprises around 100K images across 9 tissue classes. Predicting AI models’ errors in medical image classification has significant real-world implications, such as reducing misdiagnoses and increasing the reliability of AI-assisted medical tools. Appendix, Tab. S4 demonstrates that our SuperMentor accurately predicts the correctness of mentee outputs with an average accuracy of 81.8%, significantly outperforming other baseline methods. This highlights that our proposed framework can offer greater value and reliability for AI error prediction compared to existing approaches in medical image classification.
@@
| SER | 50.6 | 50.1 | 50.1 | 51.1 | 50.4 | 50.0 | 51.4 | 50.2 | 49.8 | 50.4 |
| MCP () | 53.0 | 59.1 | 53.0 | 53.2 | 50.6 | 53.8 | 53.3 | 53.2 | 54.8 | 53.8 |
| MCP () | 62.8 | 70.5 | 64.6 | 64.7 | 50.2 | 65.2 | 64.0 | 63.2 | 65.2 | 63.4 |
| MCP () | 72.5 | 78.5 | 72.1 | 75.0 | 48.8 | 77.0 | 73.8 | 76.0 | 75.2 | 72.1 |
| CPE () | 76.2 | 75.2 | 80.5 | 77.7 | 44.7 | 76.0 | 75.1 | 76.6 | 75.1 | 73.0 |
| CPE () | 75.5 | 79.5 | 76.4 | 76.2 | 48.2 | 77.8 | 76.2 | 77.6 | 76.9 | 73.8 |
| CPE () | 64.9 | 74.7 | 66.9 | 66.4 | 51.6 | 69.2 | 65.8 | 68.4 | 67.0 | 66.1 |
| DTC () | 78.9 | 66.2 | 77.8 | 80.8 | 67.1 | 79.2 | 78.0 | 78.3 | 79.1 | 76.1 |
| DTC () | 70.6 | 47.8 | 69.9 | 62.9 | 52.6 | 60.4 | 65.1 | 63.7 | 60.0 | 61.4 |
| DTC () | 51.8 | 49.5 | 52.8 | 50.3 | 50.0 | 50.8 | 51.6 | 51.8 | 50.8 | 51.0 |
| ConfidNet (Corbière et al., 2019) | 85.1 | 61.3 | 86.2 | 74.0 | 58.4 | 79.3 | 82.9 | 81.8 | 79.4 | 76.5 |
| TrustScore (Jiang et al., 2018) | 73.3 | 60.0 | 73.4 | 62.2 | 62.4 | 71.0 | 72.8 | 71.4 | 71.5 | 68.6 |
| SSL (Luo et al., 2021) | 85.3 | 58.2 | 87.8 | 75.5 | 55.4 | 77.6 | 82.8 | 79.7 | 78.0 | 75.6 |
| SuperMentor (ours) | 88.5 | 64.1 | 89.2 | 84.1 | 56.9 | 88.2 | 88.7 | 88.3 | 88.3 | 81.8 |
@@ Appendix I SuperMentor performance on additional OOD domains on ImageNet dataset
As mentioned in Sec. 5, we included experiments to evaluate the SuperMentor in more diversified OOD domains on ImageNet datasets, including ImageNet9-MIXED-RAND (IN9-MR) (Xiao et al., 2021), ImageNet9-MIXED-SAME (IN9-MS) (Xiao et al., 2021), ImageNet9-MIXED-NEXT (IN9-MN) (Xiao et al., 2021), ImageNet-R (IN-R) (Hendrycks et al., 2021a) and ImageNet-Sketch (IN-S) (Wang et al., 2019). Specifically, the MIXED-RAND, MIXED-SAME, and MIXED-NEXT datasets are derived from 9 classes in ImageNet and contain varying amounts of background and foreground signals. These datasets aim to demonstrate that models often classify objects based on background cues (often spurious features), rather than the objects themselves. Specifically, MIXED-SAME, MIXED-RAND, and MIXED-NEXT represent images with random backgrounds from the same class, random backgrounds from a random class, and random backgrounds from the next class, respectively. The ImageNet-R dataset comprises images featuring artistic representations of objects, such as cartoons, community-generated art, and graffiti renditions. The ImageNet-Sketch dataset consists of sketch-like images that match the ImageNet validation set in both categories and scale.
@@
To achieve a more thorough evaluation of our SuperMentor (see Sec. 4.5) in OOD domains, we assess its performance on these additional OOD datasets. It is worth noting that the SuperMentor learns from the mentee’s errors on adversarial ImageNet images generated by the PIFGSM attack only. In other words, the SuperMentor was NOT trained on any of these additional ImageNet-related OOD datasets. Yet, the SuperMentor achieves average accuracies of 70.9%, 71.0%, 69.6%, 62.2%, and 62.6% on the IN9-MR, IN9-MS, IN9-MN, IN-R, and IN-S datasets respectively, surpassing the 50% chance level. Despite that our Supermentor model is a simple ViT, it is still quite remarkable to achieve above-chance error prediction performance. We hope our work inspires researchers to explore more sophisticated AI mentors, incorporating advanced architectures or loss functions, to further enhance the ability to predict the correctness of another AI model’s outputs.
@@
Appendix J Detailed performance of mentors across various error sources
@@
As mentioned in the caption of Fig. 3, the detailed results of mentors across various error sources for the C10, C100, IN datasets are shown in Appendix, Fig. S4, Fig. S5 and Fig. S6 respectively.